Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O4b)
menzione - 6:44 Thursday 24 October 2024 (65381) Print this report
Operator Report - Night shift

ITF found relocking in LOW_NOISE_2 and in PREPARE_SCIENCE mode. ITF reached LN3_SQZ at 21:16 UTC. But after few minutes unlocked.
ITF is not working well. It unlocks continuously, for unknown reasons, at various steps for the whole shift.
I performed the usual cross-alignments in ACQUIRE_DRMI, and tried to lock step by step but it's not working anymore.
I decided to avoid calls to the experts OnCall becouseĀ all other ITF's was down, so, no risk to loose events.

Other - ALS
The DMS is reporting a red flag WE_ALS_Laser ALS_WEB_PSTAB_TOT_CLIP (plot1). The same issue occurred the 27/09/2023. Experts informed via mail.

Guard tour (UTC):
21:00 - 23:28
23:00 - 23:29
01:00 - 01:28
03:03 - 03:31

Sub-system reports

DET
03:30, 03:46 UTC - B1p QD1 galvo loop stuck open. Properly closed via VPM.

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×